7cosx-24sinx+5 find maximum value and minimum value

Grade:11

1 Answers

Arun
6 years ago
If we see any function of the form
a sinx + b cos x or a sinx- b cosx
Its maximum value is
\sqrt(a²+b²)
& its minimum value is
\sqrt(a²+b²)
Hence maximum value of the above function is
\sqrt(24²+ 7²)  +5 = 25 +5 = 30
& Its minimum value is
-25 +5 = -20